OpenStudy (anonymous):

Differentiate y= cos2x / x

OpenStudy (lalaly):

use the quotient rule

OpenStudy (anonymous):

i did.. but get some weird answers

OpenStudy (lalaly):

lol ok let me try

OpenStudy (anonymous):

ok thx. i first converted cos2x to cos^2 x - sin^2 x , and then i did the quotient

OpenStudy (anonymous):

i see it as math procesing error..let me open the browser again

OpenStudy (lalaly):

\[\frac{d}{dx}(\frac{\cos2x}{x})=\frac{-2xsin(2x)-\cos(2x)}{x^2}\]

OpenStudy (anonymous):

ah..then when am i suppose to conver it.... i always dont know when to convert it to that form

OpenStudy (lalaly):

\[=-\frac{2xsin2x+\cos2x}{x^2}\]

OpenStudy (anonymous):

ohh, ok thanks for ur help!

OpenStudy (lalaly):

:)